# Agentic Structured Graph Traversal for Root Cause Analysis of Code-related Incidents in Cloud Applications

### v1

@@ IV-C Phase 3: RCA Decision-making
In Phase 3 (Figure˜4), given the investigation queue $Q$ and the initial investigation narrative $I_{0}$ from Phase 2, Praxis iteratively dequeues and investigates an entity $e$ from the head of $Q$ and determines its role (a root cause vs. a symptom) in the cloud incident.
For each dequeued entity $e_{i}\in Q$ (designated as the focal entity) at iteration $i$, Praxis first constructs the observability context $c_{i}$ and retrieves the corresponding PDG $G_{P}$. It then retrieves the corresponding PDG $G_{P}$. Crucially, Praxis then employs an LLM-driven PDG traversal that traverses the entity ($e_{i}$)’s PDG ($G_{P}$) to construct an incident-centric program context $C_{P}$ that explains the observed error symptoms from the perspective of entity code context. Given the observability context $c_{i}$ and program context $C_{P}$ of the focal entity, Praxis uses the LLM to determines its role in the cloud incident as one of Primary Failure (the root cause), Symptom Only (a symptom), or Unrelated. After finishing the investigation of the current entity, Praxis moves on to investigate the other LLM-selected microservices and the dependees according to the SDG $G_{C}$ by adding them to the investigation queue $Q$. Praxis repeats this analysis on all entities in $Q$ until $Q$ is empty.
@@
Conditioned on the current state $s_{j}$, the LLM $\mathcal{M}$ produces a code-insight ($\iota_{j}$) and the graph-traversal action ($a_{j}$) pair, i.e., $\left(a_{j},\iota_{j}\right)=\mathcal{M}\left(\psi_{\tau}(s_{j})\right)$, where $\mathcal{M}$ is the LLM executing the reasoning and policy selection. $\psi_{\tau}$ is the prompt composer that encodes the current state $s_{j}$ into a structured form parametrized by a template $\tau$. In particular, $b_{j}$ and $B_{related}$’s corresponding code snippets and relations are fetched by $\psi_{\tau}$. $a_{j}\in\mathcal{T}$ is the next chosen action (mapped to our tools), and $\iota_{j}$ is the natural-language insight derived as the outcome of the current state/action. Notice that the LLM only consumes code context local to hammock blocks ($b_{j}$, $B_{related}$) for each traversal step, instead of large code files or codebases; this reduces reasoning context space and allows the LLM to focus on concise, accurate context relevant to the incident, as indicated by $c_{i}$.
The agent’s traversal action is implemented with a set of PDG traversal tools $\mathcal{T}$. Each tool in $\mathcal{T}$ corresponds to a concrete, admissible traversal action on the PDG $G_{P}$Table˜I provides a detailed summary of the above tools. When an inadmissible action (e.g., a non-existent block) is rejected by $\mathcal{T}$, Praxis prompts the LLM to regenerate a traversal action, subject to a predefined retry limit. Each action invocation updates the traversal history ($H_{i+1}$) and fetches the next node ($b_{j+1}\in G_{P}$) deterministically based on $a_{j}$ via $\mathcal{T}$ on $G_{P}$: $b_{j+1}=\mathcal{T}(G_{P};a_{j}),\;H_{j+1}=H_{j}\cup\{b_{j},\iota_{j},a_{j}\}$; we can then construct $s_{j+1}$ accordingly for the next iteration.
Figure˜5 illustrates a traversal step wherein the LLM selects the Relate action to focus on Block 1 (get_product_list function) to further investigate the observed AttributeError, as the LLM reasons that the observed AttributeError is caught within the current focal block (try-catch statement) and likely originated in get_product_list.
@@
(c) Entity role judgment and queue update. The final stage in this phase is entity role judgment and investigation queue update, in which Praxis provides a diagnostic judgment on the entity $e_{i}$’s role in the incident and an admissible explanation with evidence, and updates the investigation queue accordingly.
@@
Investigation queue update. The entity investigation queue $Q$ is then expanded to include new entities that are: (a) dependees of the current focal entity, and/or (b) suggested by the LLM based on the focal entity’s RCA decision. For (a), Praxis enqueues all unvisited dependees (i.e., cloud entities on which $e_{i}$ depends) according to the SDG $G_{C}$, denoted as $Q_{G_{C}}=G_{C}.\texttt{getDependees}(e_{i})$. For (b), Praxis prompts the LLM $\mathcal{M}$ using the entity-queuing prompt $\psi_{Q}$ to identify up to $k$ additional entities [^6] We set $k=3$ in our implementation; however, we did not observe the LLM selecting more than two entities. from $G_{C}$. This selection is based on the judgment $J$, the observability context $c_{i}$, the program context $C_{P}$, and the nodes in the SDG $G_{C}$, such that $Q_{\mathcal{M}_{Q}}=\mathcal{M}(\psi_{Q}(J,c_{i},C_{P},G_{C}))$. Finally, the queue is updated: $Q=Q\cup Q_{\mathcal{M}_{Q}}\cup Q_{G_{C}}$. Mechanism (b) is crucial as it allows the LLM to investigate entities in $G_{C}$ that surface from the program context $C_{P}$ despite having partial or missing observability data (due to insufficient coverage or the incident itself). Our evaluation confirms that this capability helps diagnose entities that do not emit observability data (e.g., a ConfigMap) or have failed silently.
@@
Praxis repeats this phase (Phase 3) for the next entity $e_{i+1}$ at the head of the updated queue $Q$. By cross-traversing between the SDG and PDG, Praxis can resolve cloud incidents that require multiple hops to root-cause the incident while maintaining a coherent RCA summary via a consistent stream of investigation narrative II–a feature that is unobtainable by simply combining independent agents [llm_rca_2023, chen2025stratus, xu2025openrca, sweagent_2024]. Critically, leveraging code-level insights to bypass gaps in the SDG, Praxis can traverse across missing links (e.g, error traces, logs) in observability data, as shown by the Cross-SDG-PDG traversal example in §III.
Moreover, by iteratively traversing the PDG node-by-node, Praxis progressively refines the RCA scope–transitioning from broader service-level observability signals to a more focused analysis of specific code regions for each focal entity. Consequently, at each step, the LLM is strictly exposed to the local context of the current graph node rather than multiple microservices or codebases/code files. This isolation of unrelated information (other node contexts) improves context efficiency, reduces token consumption, and mitigates potential context overflow, allowing Praxis to scale effectively to large cloud applications and codebases.
@@ IV-D Phase 4: Final RCA Summary
Figure˜6 presents an example final RCA report. Praxis correctly identifies the Recommendation service as the root cause, specifically attributing the failure to a schema mismatch (accessing non-existent fields),and suggests remediation. It also accurately classifies Frontend and Frontend-proxy as symptoms-only and Product-catalog as unrelated.
@@ V Experiment Setup
This section first clarifies the problem and evaluation scope, and then presents the cloud incident scenarios and implementations, baselines, and metrics used for evaluation.
@@ V-A Problem and Evaluation Scope
We evaluate Praxis on application-level incidents, focusing on microservice codebases for which source code is available for analysis. Our approach is beneficial even under partial code availability, as reasoning over available code improves diagnostic accuracy (Section˜VI); when source code is unavailable, Praxis falls back to using only observability and incident contexts for RCA. Gains are naturally limited when no code is available (Section˜VI-B). Praxis’s scope is confined to incident diagnosis within the investigation/diagnosis phase of the cloud incident lifecycle [google_sre_2016, google_cloud_incident_lifecycle].
@@ V-B Cloud Incident Evaluation Scenarios
We developed the Code-Cloud-RCA Benchmark comprising 30 cloud incidents using [opentelemetry_demo], same cloud application used successfully in [chen2025aiopslab], supported by a standard monitoring stack [jaeger_2025, prometheus_2025, clickhouse_2025, opentelemetry_overview, datadog_2025]. These scenarios are distilled from root-cause types commonly reported in production studies [cbs_database, k8saf, hotos_19, gunawi2014cloudbug, ghosh_2022_socc, yan2025empirical] and real-world incidents [k8saf, marcomutiny2024]. Table II enumerates the scenario templates, fault mechanisms, expected symptoms, and reference to real-world incidents; the rightmost column (#) records the number of scenarios per template. As with any benchmark, comprehensive coverage cannot be guaranteed, and by providing this benchmark, we invite the broader community to contribute to and expand its scope. Scenarios driven by the same template vary in timing (startup vs. serving), observability (stack traces vs. error logs vs. error codes), failure modes (stalls vs. rapid retries), or configurations, and are annotated with ground-truth root-cause microservice and reasoning. We will submit these artifacts for evaluation and open-source the benchmark.
@@
  Data schema mismatch/incompatibility between services Access a non-existent field (products_list) in a protobuf response from another microservice. Unhandled exception; 5xx bursts; request failures. (Monzo 2017 outage [monzo_2017_outage]; report [hotos_19, ghosh_2022_socc, k8saf]). 4
Improper handling of external dependency failures Upstream DB/service hangs or returns empty results; dependent service stalls without timeout or enters rapid retries. Latency spikes; retry storms; high error rate. (Discord 2023 outage [discord2023outage]; report [cbs_database, gunawi2014cloudbug, hotos_19, ghosh_2022_socc]). 6
@@
Resource-related fault Chaos Mesh: CPU stress, node-level network outage, JVM corruption. Saturation; packet loss; node/pod failures. (Multiple cases in [k8saf]) 3
@@ V-C Baselines
    ReAct-style SRE-Agent [jha2025itbench] that uses standard SRE tools to access Kubernetes API, traces, metrics, events, and microservices application logs.
@@ V-D Evaluation Metrics
  • Root cause reasoning (RCR) Pass@1 (%)
    : Pass@1 accuracy of the agent’s RCA reasoning (explanation by identifying statement(s), function(s), or configuration(s) at fault) compare against ground truth, following Pass@K [chen2021evaluating] with $K=1$.
  • Root cause identification (RCI) Pass@1 (%)
    : The percentage of instances where the identified faulty microservice(s) exactly match the ground truth.
We also report RCA efficiency via MTTD (mean time to diagnosis) and ATC (average token consumption).
@@ V-E Praxis Implementation
Praxis is implemented in Python 3.12 using the LangGraph agentic framework [langgraph_2025]. Praxis supports PDG construction for microservices written in Python and Java, a scope we chose based on current toolchain availability and the popularity of these languages [tiobe2025index]. For evaluation purposes, all faults implemented reside in codebases available for analysis and supported by Praxis. We will submit our artifacts for evaluation and open-source the implementation.
@@ VI Evaluation Results
  gpt-oss-120b 2.68 $\pm$ 1.32 8.72 $\pm$ 2.31
deepseek-r1 5.48 $\pm$ 1.88 20.44 $\pm$ 3.34
o4-mini 9.86 $\pm$ 2.50 22.54 $\pm$ 3.50
mistral-medium-3.1 14.62 $\pm$ 3.10 28.46 $\pm$ 3.96
gpt-5 19.71 $\pm$ 3.40 35.77 $\pm$ 4.10
gpt-5-codex 5.37 $\pm$ 1.85 13.42 $\pm$ 2.79
                           SRE-Agent + Code Tools (CT)
  gpt-oss-120b 2.01 $\pm$ 1.15 6.71 $\pm$ 2.01
deepseek-r1 4.68 $\pm$ 2.64 17.18 $\pm$ 4.72
o4-mini 7.63 $\pm$ 2.21 24.31 $\pm$ 3.57
mistral-medium-3.1 16.15 $\pm$ 3.22 27.69 $\pm$ 3.93
gpt-5 25.55 $\pm$ 3.73 36.50 $\pm$ 4.11
gpt-5-codex 5.63 $\pm$ 1.93 10.56 $\pm$ 2.58
@@
  • Metrics: (i) RCR Pass@1: Pass@1 % score for root cause reasoning; (ii) RCI Pass@1: root cause identification accuracy. Best-performing agent+LLM model is shown in bold, and the second-best is underlined. Higher is better.
@@ VI-A Baseline Comparison (RQ1)
For RQ-1, we evaluated Praxis by comparing it against the current state of the art, i.e., the SRE-Agent [jha2025itbench], a ReAct-style RCA agent with default tools, and an extended SRE-Agent with code retrieval and summarization tools (denoted by SRE-Agent+CT) to study whether code context can improve on the baseline performance.
@@
Evaluation setup. We evaluated the baselines and Praxis across six different LLMs: o4-mini, gpt-5, gpt-5-codex, mistral-medium-3.1, deepseek-r1, and gpt-oss-120b. For statistical robustness, we repeated each $\langle\text{LLM},\ \text{scenario}\rangle$ experiment five times with different random seeds, totaling 2,700 trajectories.
@@
Overall results. As shown in Table˜III, Praxis consistently performed better than SRE-Agent and SRE-Agent+CT across the LLMs that were evaluated, both on root-cause reasoning (RCR) accuracy and root-cause identification (RCI) accuracy. Praxis’s RCA reasoning accuracy correlates strongly with root-cause identification accuracy, reflecting how human SREs rely on accurate fault pinpointing for effective root-cause analysis.
For ReAct-style SRE agent baselines, RCA accuracy is tightly coupled to tool-calling: models with stronger tool use (mistral-medium-3.1, gpt-5) achieve higher RCA accuracy, whereas strong reasoning models not tuned for tool usage (e.g., o4-mini, deepseek-r1) underperform due to inconsistent tool calls. These inconsistencies cause them to invoke incorrect tools, pass incorrect parameters, or skip necessary tools, leading investigations to end with symptom-level explanations (e.g., error traces) rather than true root causes. The gap between gpt-5 and gpt-5-codex within the same model family manifests this issue: gpt-5-codex, tuned primarily for code-centric tasks and tools, attains lower RCA accuracy than gpt-5, which is tuned for more general tool-calling and thus generalizes better to both SRE and code tools required for RCA.
For Praxis, RCA accuracy correlates more closely with the model’s reasoning capabilities: models with stronger reasoning (gpt-5, gpt-5-codex, o4-mini) exhibit higher RCA accuracy overall. Moreover, we speculate that the stronger performance of gpt-5-codex compared to gpt-5 stems from its superior code-understanding capabilities (as demonstrated on [jimenez2024swebench]), which better support correlating code context and program behavior with the incident. The low accuracy of mistral-medium-3.1 is primarily due to its failure to select appropriate initial candidates and its inability to follow output-format instructions, which frequently led to output parsing errors.
@@
Why did baselines underperform? SRE-Agent underperformed because it lacks code visibility, which is essential for diagnosing code-related scenarios. Adding code tools (SRE-Agent+CT) improves accuracy with gpt-5 but yields only marginal gains for other LLMs. This is not because code tools are unhelpful, but because the baseline agents’ investigation decision-making remains unstructured and unguided. Baseline ReAct-style agents make diagnosis decisions (as additional tool-calls or termination) primarily from the most recent tool-call states and observations, without an explicit representation of or planning over the underlying dependency graphs, making their RCA accuracy heavily dependent on the model’s tool-calling capabilities. As a result, even with both code and SRE tools, models not tuned for general tool use (e.g., o4-mini, gpt-5-codex) tend to exhibit two failure modes: (1) premature closure: once local evidence appears to explain the alerts (e.g., an HTTP 500 handler or related error traces), they terminate the investigation without diving deeper; and (2) tool-usage fragility: invoking incorrect tools or supplying incorrect parameters [winston_tool-augmentedLLMs_2025, xiong-etal-2025-butterfly]. Additionally, without explicit awareness or external imposition of the microservice or program dependencies, these baselines had neither a mechanism nor an obligation to traverse the multi-hop dependency chain (which in our scenarios often involved 7–10 hops across different microservices and their hammock blocks) in which the true root-causes were hinted or resided.
@@
Notably, even with explicit instructions in the prompt to invoke code tools, SRE-Agent+CT with gpt-5, the best-performing model, invoked code tools in only 32% of runs where code analysis was required for RCA. This under-use of required tools led to lower RCA accuracy and inconsistent execution trajectories, reinforcing the baseline agent failure modes described above. In contrast, Praxis treats service and code-level dependencies as first-class as it constrains the LLM’s reasoning onto the dependency paths in PDG and systematically expands each graph node’s neighborhood with admissible graph traversal actions, enforcing iterative context gathering and reasoning reconciliation. This graph-aware reasoning prevents early stopping on symptomatic context and steers the investigation to the upstream root cause in code, achieving in-depth RCA.

### v2

@@ IV-C Phase 3: RCA Decision-making
In Phase 3 (Figure˜4), given the investigation queue $Q$ and the initial investigation narrative $I_{0}$ from Phase 2, Praxis iteratively dequeues and investigates an entity $e$ from the head of $Q$ and determines its role: a root cause, a symptom, or unrelated to the cloud incident.
For each dequeued entity $e_{i}\in Q$ (designated as the focal entity) at iteration $i$, Praxis first constructs the observability context $c_{i}$ and retrieves the corresponding PDG $G_{P}$. Crucially, Praxis then employs an LLM-driven PDG traversal that traverses the entity’s ($e_{i}$’s) PDG ($G_{P}$) to construct an incident-centric program context $C_{P}$ that explains the observed error symptoms from the perspective of entity code context. Given the observability context $c_{i}$ and program context $C_{P}$ of the focal entity, Praxis uses the LLM to determine its role in the cloud incident as one of Primary Failure (the root cause), Symptom Only (a symptom), or Unrelated. After finishing the investigation of the current entity, Praxis moves on to investigate the other LLM-selected microservices and the dependees according to the SDG $G_{C}$ by adding them to the investigation queue $Q$. Praxis repeats this analysis on all entities in $Q$ until $Q$ is empty.
@@
Conditioned on the current state $s_{j}$, the LLM $\mathcal{M}$ produces a code-insight ($\iota_{j}$) and the graph-traversal action ($a_{j}$) pair, i.e., $\left(a_{j},\iota_{j}\right)=\mathcal{M}\left(\psi_{\tau}(s_{j})\right)$, where $\mathcal{M}$ is the LLM executing the reasoning and policy selection. $\psi_{\tau}$ is the prompt composer that encodes the current state $s_{j}$ into a structured form parametrized by a template $\tau$. In particular, $b_{j}$ and $B_{related}$’s corresponding code snippets and relations are fetched by $\psi_{\tau}$. $a_{j}\in\mathcal{T}$ is the next chosen action (mapped to our tools), and $\iota_{j}$ is the natural-language insight derived as the outcome of the current state/action. Notice that the LLM only consumes code context local to hammock blocks ($b_{j}$, $B_{related}$) for each traversal step, instead of large code files or codebases; this reduces the reasoning context space and allows the LLM to focus on concise, accurate context relevant to the incident, as indicated by $c_{i}$.
The agent’s traversal action is implemented with a set of PDG traversal tools $\mathcal{T}$. Each tool in $\mathcal{T}$ corresponds to a concrete, admissible traversal action on the PDG $G_{P}$Table˜I provides a detailed summary of the above tools. When an inadmissible action (e.g., a nonexistent block) is rejected by $\mathcal{T}$, Praxis prompts the LLM to regenerate a traversal action, subject to a predefined retry limit. Each action invocation updates the traversal history ($H_{i+1}$) and fetches the next node ($b_{j+1}\in G_{P}$) deterministically based on $a_{j}$ via $\mathcal{T}$ on $G_{P}$: $b_{j+1}=\mathcal{T}(G_{P};a_{j}),\;H_{j+1}=H_{j}\cup\{b_{j},\iota_{j},a_{j}\}$; we can then construct $s_{j+1}$ accordingly for the next iteration.
Figure˜5 illustrates a PDG traversal step in which the LLM selects the Relate action to focus on Block 1 (the get_product_list function) for further investigation of the observed AttributeError. The LLM reasons that although the error is caught in the current focal block (the try-catch statement), it likely arises during execution of get_product_list.
@@
(c) Entity role judgment and queue update. The final stage in this phase is entity role judgment and an investigation queue update, in which Praxis provides a diagnostic judgment on the entity $e_{i}$’s role in the incident and an admissible explanation with evidence, and updates the investigation queue accordingly.
@@
Investigation queue update. The entity investigation queue $Q$ is then expanded to include new entities that are (1) dependees of the current focal entity, and/or (2) suggested by the LLM based on the focal entity’s RCA decision. For (1), Praxis enqueues all unvisited dependees of $e_{i}$ according to the SDG $G_{C}$, denoted by $Q_{G_{C}}=G_{C}.\texttt{getDependees}(e_{i})$, where each dependee is a cloud entity on which $e_{i}$ depends. For (2), Praxis prompts the LLM $\mathcal{M}$ using the entity-queuing prompt $\psi_{Q}$ to identify up to $k$ additional entities [^5] We set $k=3$ in our implementation; however, across all evaluation runs, the LLM selected at most two entities. from $G_{C}$. This selection is based on the judgment $J$, the observability context $c_{i}$, the program context $C_{P}$, and the nodes in the SDG $G_{C}$, such that $Q_{\mathcal{M}_{Q}}=\mathcal{M}(\psi_{Q}(J,c_{i},C_{P},G_{C}))$. Finally, the queue is updated: $Q=Q\cup Q_{\mathcal{M}_{Q}}\cup Q_{G_{C}}$. Mechanism (2) is crucial as it allows the LLM to investigate entities in $G_{C}$ that surface from the program context $C_{P}$ despite having partial or missing observability data (due to insufficient coverage or the incident itself). Our evaluation confirms that this capability helps diagnose entities that do not emit observability data (e.g., a ConfigMap) or have failed silently.
@@
Praxis repeats this phase (Phase 3) for the next entity $e_{i+1}$ at the head of the updated queue $Q$. By cross-traversing between the SDG and PDG, Praxis can resolve cloud incidents that require multiple hops to root-cause the incident while maintaining a coherent RCA summary via a consistent stream of investigation narrative $I$, a feature that is unobtainable by simply combining independent agents [llm_rca_2023, chen2025stratus, xu2025openrca, sweagent_2024]. Critically, leveraging code-level insights to bypass gaps in the SDG, Praxis can traverse across missing links (e.g, error traces, logs) in observability data, as shown by the Cross-SDG-PDG traversal example in §III.
Moreover, by iteratively traversing the PDG node-by-node, Praxis progressively refines the RCA scope, transitioning from broader service-level observability signals to a more focused analysis of specific code regions for each focal entity. Consequently, at each step, the LLM is strictly exposed to the local context of the current graph node rather than multiple microservices or codebases/code files. This isolation of unrelated information (other node contexts) improves context efficiency, reduces token consumption, and mitigates potential context overflow, allowing Praxis to scale effectively to large cloud applications and codebases.
@@ IV-D Phase 4: Final RCA Summary
Figure˜6 presents an example final RCA report. Praxis correctly identifies the Recommendation service as the root cause, specifically attributing the failure to a schema mismatch (accessing nonexistent fields), and suggests remediation. It also accurately classifies Frontend and Frontend-proxy as Symptom Only and Product-catalog as Unrelated.
@@ V Experiment Setup
This section first clarifies the problem and evaluation scope, and then presents the cloud incident scenarios, baselines, and metrics used for evaluation.
@@ V-A Problem and Evaluation Scope
We evaluate Praxis on application-level incidents, focusing on microservice codebases for which source code is available for analysis. Our approach is beneficial even under partial code availability, as reasoning over available code improves diagnostic accuracy (see §VI-A); when source code is unavailable, Praxis falls back to using only observability and incident contexts for RCA. Gains are naturally limited when no code is available (see §VI-B). Praxis’s scope is confined to incident diagnosis within the investigation/diagnosis phase of the cloud incident lifecycle [google_sre_2016, google_cloud_incident_lifecycle].
@@ V-B Cloud Incident Evaluation Scenarios
We developed the Code-Cloud-RCA Benchmark comprising 30 cloud incidents using [opentelemetry_demo], the same cloud application used in [chen2025aiopslab], supported by a standard monitoring stack [jaeger_2025, prometheus_2025, clickhouse_2025, opentelemetry_overview, datadog_2025]. These scenarios are distilled from root-cause types commonly reported in production studies [cbs_database, k8saf, hotos_19, gunawi2014cloudbug, ghosh_2022_socc, yan2025empirical] and real-world incidents [k8saf, marcomutiny2024]. Table II enumerates the scenario templates, fault mechanisms, expected symptoms, and reference to real-world incidents; the rightmost column (#) records the number of scenarios per template. As with any benchmark, comprehensive coverage cannot be guaranteed, and by providing this benchmark, we invite the broader community to contribute to and expand its scope. Scenarios driven by the same template vary in timing (startup vs. serving), observability (stack traces vs. error logs vs. error codes), failure modes (stalls vs. rapid retries), or configurations, and are annotated with ground-truth root-cause microservices and reasoning.
@@
  Data schema mismatch/incompatibility between services Access a nonexistent field (products_list) in a protobuf response from another microservice. Unhandled exception; 5xx bursts; request failures. (Monzo 2017 outage [monzo_2017_outage]; report [hotos_19, ghosh_2022_socc, k8saf]) 4
Improper handling of external dependency failures Upstream DB/service hangs or returns empty results; dependent service stalls without timeout or enters rapid retries. Latency spikes; retry storms; high error rate. (Discord 2023 outage [discord2023outage]; report [cbs_database, gunawi2014cloudbug, hotos_19, ghosh_2022_socc]) 6
@@
Resource- & Infrastructure-related fault Chaos Mesh: CPU stress, node-level network outage, JVM corruption. Saturation; packet loss; node/pod failures. (Multiple cases in [k8saf]) 3
@@ V-C Baselines
    ReAct-style SRE-Agent [jha2025itbench] that uses standard SRE tools to access the Kubernetes API, traces, metrics, events, and microservices application logs.
@@ V-D Evaluation Metrics
  • Root cause identification (RCI) Pass@1 (%):
    The percentage of instances for which the identified faulty microservice(s) exactly match the ground truth. In other words, RCI measures whether the agent correctly localizes the root cause to the responsible microservice(s) or associated resource(s).
  • Root cause reasoning (RCR) Pass@1 (%):
    Pass@1 accuracy of the agent’s RCA reasoning, measured by whether its explanation correctly identifies the faulty statement(s), function(s), or configuration(s) relative to the ground truth, following Pass@K [chen2021evaluating] with $K=1$. In other words, RCR measures whether the agent pinpoints the underlying fault (e.g., a faulty code-site or configuration) and correctly explains (reasons about) how it led to the incident and the observed symptoms.
We also report RCA efficiency via MTTD (mean time to diagnosis) and ATC (average token consumption). See §VI-C for a detailed explanation of MTTD and ATC.
@@ V-E Praxis Implementation
Praxis is implemented in Python 3.12 using the LangGraph agentic framework [langgraph_2025]. Praxis supports PDG construction for microservices written in Python and Java, a scope we chose based on current toolchain availability and the popularity of these languages [tiobe2025index]. For evaluation purposes, all implemented faults reside in, or are associated with, codebases available for analysis and supported by Praxis. We will submit our artifacts for evaluation and open-source the implementation.
@@ VI Evaluation Results
  gpt-oss-120b 0.00 $\pm$ 0.00 1.53 $\pm$ 1.07
deepseek-r1 0.00 $\pm$ 0.0 9.79 $\pm$ 2.49
o4-mini 3.42 $\pm$ 1.50 11.64 $\pm$ 2.65
mistral-medium-3.1 9.65 $\pm$ 2.77 21.93 $\pm$ 3.88
gpt-5-codex 0.00 $\pm$ 0.00 11.43 $\pm$ 3.80
                           SRE-Agent w/ Code Tools (CT)
  gpt-oss-120b 0.68 $\pm$ 0.69 1.38 $\pm$ 0.97
deepseek-r1 0.00 $\pm$ 0.0 11.41 $\pm$ 2.60
o4-mini 2.68 $\pm$ 1.32 14.77 $\pm$ 2.91
mistral-medium-3.1 10.07 $\pm$ 2.55 20.86 $\pm$ 3.45
gpt-5-codex 0.00 $\pm$ 0.00 7.96 $\pm$ 2.54
@@
  • Metrics: (1) RCR Pass@1: Pass@1 % score for root cause reasoning. (2) RCI Pass@1: root cause identification accuracy. Best-performing agent+LLM model is shown in bold, and the second-best is underlined. Higher is better.
@@ VI-A Baseline Comparison (RQ1)
For RQ-1, we evaluated Praxis by comparing it against the current state of the art, i.e., the SRE-Agent [jha2025itbench], a ReAct-style RCA agent with default tools, and an extended SRE-Agent with code retrieval and summarization tools (which we call SRE-Agent+CT) to study whether code context can improve on the baseline performance.
@@
Evaluation setup. We evaluated the agent across five different LLMs: o4-mini, gpt-5-codex, mistral-medium-3.1, deepseek-r1, and gpt-oss-120b. For statistical robustness, we repeated each $\langle\text{LLM},\ \text{scenario}\rangle$ experiment five times with distinct random seeds, totaling 2,250 trajectories.
@@
Overall results. As shown in Table˜III, Praxis consistently performed better than SRE-Agent and SRE-Agent+CT across all LLMs that were evaluated, both on root-cause reasoning (RCR) accuracy and root-cause identification (RCI) accuracy. Praxis’s RCA reasoning accuracy correlates strongly with root-cause identification accuracy, reflecting how human SREs rely on accurate fault pinpointing for effective root-cause analysis. Reasoning models (gpt-5-codex, o4-mini) exhibited better RCA accuracies overall. Moreover, the strong performance of gpt-5-codex likely stemmed from its superior code-understanding capabilities (demonstrated in [jimenez2024swebench]), enabling it to better correlate the code context and program behavior with the incident. The low accuracy of mistral-medium-3.1 was due to its failure to select the correct initial candidates and its inability to follow output format instructions, which resulted in parsing errors.
@@
Why did baselines underperform? SRE-Agent underperformed because it lacks code visibility, which is essential for diagnosing code-related scenarios. Adding code tools (SRE-Agent+CT) yielded only marginal gains, not because code is unhelpful, but because the search policy remained unguided. Even with code and SRE tools, gpt-5-codex and similar models exhibited premature closure: once local evidence seemed to explain alerts (e.g., an HTTP 500 handler or nearby error traces), they terminated the investigation. Without explicit awareness or external imposition of the microservice or program dependencies, these baselines had neither a mechanism for traversing, nor an obligation to traverse the multi-hop dependency chain (which in our scenarios often involved 7–10 hops across different microservices and their hammock blocks) in which the true faults either resided or were indicated. Notably, despite explicit instructions to invoke code tools in the prompt, SRE-Agent+CT invoked code tools for only 46% of runs in which code analysis was required for RCA, resulting in low RCA accuracy and inconsistent execution trajectories and highlighting SRE-Agent+CT’s fragility in agent tool-calling [winston_tool-augmentedLLMs_2025, xiong-etal-2025-butterfly].
@@
In contrast, Praxis treats service and code-level dependencies as first-class as it constrains the LLM’s reasoning onto the dependency paths in PDG and systematically expands each graph node’s neighborhood with admissible graph traversal actions, enforcing iterative context gathering and reasoning reconciliation. This graph-aware reasoning prevents early stopping on symptomatic context and steers the investigation to the upstream root cause in code, achieving in-depth RCA.